Corporate triple A bond interest rates for 12 consecutive months are as follows:

9.6 9.2 9.3 9.7 9.8 9.7 9.7 10.6 9.9 9.6 9.4 9.7

If required, round your answer to two decimal places

(b) Develop three-month and four-month moving averages for this time series.
If required, round your answers to two decimal places.

Week

Sales
3 Month
Moving Average
4 Month
Moving Average
1 9.6
2 9.2
3 9.3
4 9.7
5 9.8
6 9.7
7 9.7
8 10.6
9 9.9
10 9.6
11 9.4
12 9.7
Enter the Mean Square Errors for the three-month and the four-month moving average forecasts. If needed, round your answers to three decimal digits.
3-month
moving average
4-month
moving average
MSE
Does the three-month or the four-month moving average provide the better forecasts based on MSE? Explain.
- Select your answer -3-month moving average4-month moving averageItem 22
- Select your answer -Moving average with the lowest MSE provides the better forecast.Moving average with the highest MSE provides the better forecast.Item 23
(c) What is the moving average forecast for the next month?
Answer #1

answer:

Develop three-month and four-month moving averages for this time series.

MSE of 3 month Moving Average = 0.155

MSE of 4 month Moving Average = 0.172

The three month moving average provide the better forecast as we can see that the MSE for 3 month is less than 4 month moving average.

c) 3 Month Moving Average forecast for next month = (9.6 + 9.7 + 9.5) / 3 = 9.6

4 Month Moving Average forecast for next month = (10.07 +10.07 +10.03 + 9.73) / 4 = 9.98